# Diffusion State-Guided Projected Gradient for Inverse Problems

### v2

@@ Abstract
Recent advancements in diffusion models have been effective in learning data priors for solving inverse problems. They leverage diffusion sampling steps for inducing a data prior while using a measurement guidance gradient at each step to impose data consistency. For general inverse problems, approximations are needed when an unconditionally trained diffusion model is used since the measurement likelihood is intractable, leading to inaccurate posterior sampling. In other words, due to their approximations, these methods fail to preserve the generation process on the data manifold defined by the diffusion prior, leading to artifacts in applications such as image restoration. To enhance the performance and robustness of diffusion models in solving inverse problems, we propose Diffusion State-Guided Projected Gradient (DiffStateGrad), which projects the measurement gradient onto a subspace that is a low-rank approximation of an intermediate state of the diffusion process. DiffStateGrad, as a module, can be added to a wide range of diffusion-based inverse solvers to improve the preservation of the diffusion process on the prior manifold and filter out artifact-inducing components. We highlight that DiffStateGrad improves the robustness of diffusion models in terms of the choice of measurement guidance step size and noise while improving the worst-case performance. Finally, we demonstrate that DiffStateGrad improves upon the state-of-the-art on linear and nonlinear image restoration inverse problems.
@@ 2 Background & Related Works
Solving inverse problems with diffusion models. Diffusion-based approaches to inverse problems seek to reconstruct the original data ${\bm{x}}_{0}$ from the measurement ${\bm{y}}=\mathcal{A}({\bm{x}}_{0})+{\bm{n}}$. In this case, the reverse SDE implements
@@
Gradient guidance incorporation.  Prior works differ from one another in two key aspects: (a) how the gradient is computed and (b) how the gradient is used to update the diffusion state. Table 2 categorizes prior works based on these characteristics. While a few approaches, such as diffusion-based MRI (Chung & Ye, 2022), compute the measurement gradient using ${\bm{x}}_{t}$, most recent literature has shifted toward using ${\bm{x}}_{0|t}$ for gradient computation. Regarding gradient incorporation, the literature is further subdivided. For instance, methods like DPS and PSLD use the measurement gradient to update the state at time $t$, whereas other approaches, such as ReSample, DAPS, and MPGD, apply the guidance to the conditional state at $0|t$ before resampling. Additionally, while the projections in MCG (Chung et al., 2022a) and MPGD (He et al., 2024) are restricted to ${\bm{x}}_{0|t}$, DiffStateGrad stands out by being applicable to both types of methods.
@@ 3 Diffusion State-Guided Projected Gradient (DiffStateGrad)
We propose a Diffusion State-Guided Projected Gradient (DiffStateGrad) to solve inverse problems. DiffStateGrad can be incorporated into a wide range of diffusion models to improve guidance-based diffusion models. Without loss of generality, we explain DiffStateGrad in the context of Latent-DPS (Chung et al., 2023) which applies the measurement guidance to ${\bm{z}}_{t}$. We note that DiffStateGrad applies to a wide range of pixel and latent diffusion-based inverse solvers (see Section 4).
@@
where $\bm{G}_{t}$ is the measurement gradient in image matrix form, and $\bm{U}_{r}$ and $\bm{V}_{r}$ contain the first $r$ left and right singular vectors, respectively (Algorithm 1). While we use the full SVD projection (i.e., combining both left and right projection), in practice, one may choose to do either left or right projection. Next, we provide mathematical intuitions (Proposition 1) on the effectiveness of DiffStateGrad in preserving ${\bm{z}}_{t}$, particularly for high-dimensional data with low-rank structure, after the MG update on the manifold $\mathcal{M}_{t}$. Finally, we note that while DiffStateGrad can significantly improve the runtime and computational efficiency of diffusion frameworks that use Adam optimizers for data consistency (Song et al., 2023a; Zhao et al., 2024), the current implementation and this paper does not explore this aspect and, instead, focuses on the property of the proposed subspace.
@@
Efficiency. DiffStateGrad introduces minimal computational overhead. We perform SVD at most once per iteration, and for latent diffusion solvers, this occurs in the latent space on $64\times 64$ matrices. By selecting a low rank based on a variance threshold, subsequent projection and reconstruction operations are performed on reduced matrices, further decreasing computational complexity. Figure 4 illustrates the runtime for the three diffusion-based methods of PSLD (Rout et al., 2023), ReSample (Song et al., 2023a), and DAPS (Zhang et al., 2024) with and without DiffStateGrad. The figure shows that the additional computational cost of incorporating DiffStateGrad is marginal, typically adding only a few seconds to the total runtime (see C.2 for further details). We note that our method is not intended to improve efficiency, but rather to enhance performance and robustness.
@@ 4 Results
This section provides extensive experimental results on the effectiveness of DiffStateGrad employed in several methods for image-based inverse problems. We show that DiffStateGrad significantly improves (1) the robustness of diffusion-based methods to the choice of measurement gradient step size and measurement noise, and (2) the overall posterior sampling performance of diffusion.
@@
Performance. We provide quantitative results in Tables 3(a), 3(b) and 4(b), and qualitative results in Figure 8. These results demonstrate the substantial improvement in the performance of DiffStateGrad-PSLD, DiffStateGrad-ReSample, DiffStateGrad-DPS, and DiffStateGrad-DAPS against their respective SOTA counterparts across a wide variety of linear and nonlinear tasks on both FFHQ and ImageNet. For example, DiffStateGrad significantly improves performance and increases reconstruction consistency in both pixel-based and latent solvers for phase retrieval (Table 3(b)). Our results also highlight the superiority of our choice of subspace compared to prior works. For example, DiffStateGrad-DPS is superior to MCG, which is the special case of DPS for manifold constrained diffusion. Furthermore, the outperformance of DiffStateGrad-DAPS against DAPS and MPGD-AE emphasizes the effectiveness of the proposed SVD-based subspace projection. We additionally show that DiffStateGrad does not impact the diversity of the posterior sampling (see Figure 17). Finally, we note that DiffStateGrad is robust to the choice of subspace rank (Figures 9 and 10). We refer the reader to Appendix B for the extensive qualitative performance of DiffStateGrad.
@@ 6 Reproducibility Statement
To ensure the reproducibility of our results, we thoroughly detail the hyperparameters used in our experiments in C.1. We also provide specific implementation and configuration details of all the baselines used in C.3C.4C.5, and C.6. Moreover, we use easily accessible pre-trained diffusion models throughout our experiments. PSLD (https://github.com/LituRout/PSLD), ReSample (https://github.com/soominkwon/resample), DPS (https://github.com/DPS2022/diffusion-posterior-sampling), and DAPS (https://github.com/zhangbingliang2019/DAPS) all have publicly available code. We will make our code available upon acceptance to the conference.
@@ Appendix A Additional Results
Table 6: SSIM comparison for linear tasks on FFHQ 256 $\times$ 256.
@@
Table 7: SSIM comparison for nonlinear tasks on FFHQ 256 $\times$ 256.

### v3

@@ Abstract
Recent advancements in diffusion models have been effective in learning data priors for solving inverse problems. They leverage diffusion sampling steps for inducing a data prior while using a measurement guidance gradient at each step to impose data consistency. For general inverse problems, approximations are needed when an unconditionally trained diffusion model is used since the measurement likelihood is intractable, leading to inaccurate posterior sampling. In other words, due to their approximations, these methods fail to preserve the generation process on the data manifold defined by the diffusion prior, leading to artifacts in applications such as image restoration. To enhance the performance and robustness of diffusion models in solving inverse problems, we propose Diffusion State-Guided Projected Gradient (DiffStateGrad), which projects the measurement gradient onto a subspace that is a low-rank approximation of an intermediate state of the diffusion process. DiffStateGrad, as a module, can be added to a wide range of diffusion-based inverse solvers to improve the preservation of the diffusion process on the prior manifold and filter out artifact-inducing components. We highlight that DiffStateGrad improves the robustness of diffusion models in terms of the choice of measurement guidance step size and noise while improving the worst-case performance. Finally, we demonstrate that DiffStateGrad improves upon the state-of-the-art on linear and nonlinear image restoration inverse problems. Our code is available at https://github.com/neuraloperator/DiffStateGrad.
@@ 2 Background & Related Works
Solving inverse problems with diffusion models. Diffusion-based approaches to inverse problems seek to reconstruct the data ${\bm{x}}_{0}$ from the measurement ${\bm{y}}=\mathcal{A}({\bm{x}}_{0})+{\bm{n}}$ via the following reverse SDE
@@
Gradient guidance incorporation. Prior works differ from one another in two key aspects: (a) how the loss for the gradient is computed and (b) how the gradient is used to update the diffusion state. Table 2 categorizes prior works based on these characteristics. While a few approaches, such as diffusion-based MRI (Chung & Ye, 2022), compute the gradient using ${\bm{x}}_{t}$, most recent literature has shifted toward using ${\bm{x}}_{0|t}$ for gradient computation. Regarding gradient incorporation, the literature is further subdivided. For instance, methods like DPS and PSLD use the measurement gradient to update the state at time $t$, whereas ReSample, DAPS, and MPGD apply the guidance to the conditional state at $0|t$ before resampling. Additionally, while the projections in MCG (Chung et al., 2022a) and MPGD (He et al., 2024) are restricted to ${\bm{x}}_{0|t}$, DiffStateGrad applies to both types of methods.
@@ 3 Diffusion State-Guided Projected Gradient (DiffStateGrad)
We propose a Diffusion State-Guided Projected Gradient (DiffStateGrad) to solve inverse problems. DiffStateGrad can be incorporated into a wide range of diffusion models to improve guidance-based diffusion models. Without loss of generality, we explain DiffStateGrad in the context of Latent-DPS (Chung et al., 2023) hich applies the measurement guidance to ${\bm{z}}_{t}$. We note that DiffStateGrad applies to a wide range of pixel and latent diffusion-based inverse solvers (see Section 4).
@@
where $\bm{G}_{t}$ is the measurement gradient in image matrix form, and $\bm{U}_{r}$ and $\bm{V}_{r}$ contain the first $r$ left and right singular vectors, respectively (Algorithm 1). While we use the full SVD projection (i.e., combining both left and right projection), in practice, one may choose to do either left or right projection. Next, we provide mathematical intuitions (Proposition 1) on the effectiveness of subspace projections in preserving ${\bm{z}}_{t}$, particularly for high-dimensional data with low-rank structure, after the MG update on the manifold $\mathcal{M}_{t}$. Finally, we note that while DiffStateGrad can significantly improve the runtime and computational efficiency of diffusion frameworks that use Adam optimizers for data consistency (Song et al., 2023a; Zhao et al., 2024), the current implementation and this paper does not explore this aspect and, instead, focuses on the property of the proposed subspace.
@@
Efficiency. DiffStateGrad introduces minimal computational overhead. We perform SVD at most once per iteration, and for latent diffusion solvers, this occurs in the latent space on $64\times 64$ matrices. By selecting a low rank based on a variance threshold, subsequent projection and reconstruction operations are performed on reduced matrices, further decreasing computational complexity. Figure 4 illustrates the runtime of PSLD (Rout et al., 2023), ReSample (Song et al., 2023a), and DAPS (Zhang et al., 2024) with and without DiffStateGrad. The figure shows that the additional computational cost of incorporating DiffStateGrad is marginal, typically adding only a few seconds to the total runtime (see C.2 for further details). We note that our method is not intended to improve efficiency, but rather to enhance performance and robustness.
@@ 4 Results
This section provides extensive experimental results on the effectiveness of DiffStateGrad for image-based inverse problems. We show that DiffStateGrad significantly improves (1) the robustness of diffusion-based methods to the choice of measurement gradient step size and measurement noise, and (2) the overall posterior sampling performance of diffusion.
@@
Performance. We provide quantitative results in Tables 3(a), 3(b) and 4(b), and qualitative results in Figure 8. These results demonstrate the substantial improvement in the performance of the methods with DiffStateGrad against their respective SOTA counterparts across a wide variety of linear and nonlinear tasks on both FFHQ and ImageNet. For example, DiffStateGrad significantly improves performance and increases reconstruction consistency in both pixel-based and latent solvers for phase retrieval (Table 3(b)). Our results also highlight the superiority of our choice of subspace compared to prior works. For example, DiffStateGrad-DPS is superior to MCG, which is the special case of DPS for manifold constrained diffusion. Furthermore, the outperformance of DiffStateGrad-DAPS against DAPS and MPGD-AE emphasizes the effectiveness of the proposed SVD-based subspace projection. We additionally show that DiffStateGrad does not impact the diversity of the posterior sampling (Figure 17). Finally, we note that DiffStateGrad is robust to the choice of subspace rank (Figures 9 and 10). We refer the reader to Appendix B for the extensive qualitative performance of DiffStateGrad.
@@ 6 Reproducibility Statement
To ensure the reproducibility of our results, we thoroughly detail the hyperparameters used in our experiments in C.1. We also provide specific implementation and configuration details of all the baselines used in C.3C.4C.5, and C.6. Moreover, we use easily accessible pre-trained diffusion models throughout our experiments. PSLD (https://github.com/LituRout/PSLD), ReSample (https://github.com/soominkwon/resample), DPS (https://github.com/DPS2022/diffusion-posterior-sampling), and DAPS (https://github.com/zhangbingliang2019/DAPS) all have publicly available code. We also make our code available at https://github.com/neuraloperator/DiffStateGrad.
@@ Appendix A Additional Results
Table 6: SSIM comparison on FFHQ 256 $\times$ 256.
(a) Linear tasks.
@@
(b) Nonlinear tasks.